# Raising the Bar: Investigating the Values of Large Language Models via Generative Evolving Testing

### v1

@@ Abstract
Warning: this paper contains model outputs exhibiting unethical information. Large Language Models (LLMs) have achieved significant breakthroughs, but their generated unethical content poses potential risks. Measuring value alignment of LLMs becomes crucial for their regulation and responsible deployment. Numerous datasets have been constructed to assess social bias, toxicity, and ethics in LLMs, but they suffer from evaluation chronoeffect, that is, as models rapidly evolve, existing data becomes leaked or undemanding, overestimating ever-developing LLMs. To tackle this problem, we propose GETA, a novel generative evolving testing approach that dynamically probes the underlying moral baselines of LLMs. Distinct from previous adaptive testing methods that rely on static datasets with limited difficulty, GETA incorporates an iteratively-updated item generator which infers each LLM’s moral boundaries and generates difficulty-tailored testing items, accurately reflecting the true alignment extent. This process theoretically learns a joint distribution of item and model response, with item difficulty and value conformity as latent variables, where the generator co-evolves with the LLM, addressing chronoeffect. We evaluate various popular LLMs with diverse capabilities and demonstrate that GETA can create difficulty-matching testing items and more accurately assess LLMs’ values, better consistent with their performance on unseen OOD and i.i.d. items, laying the groundwork for future evaluation paradigms.
**footnotetext: Corresponding authors. This work was done during the first author’s internship at MSRA.
@@ 1 Introduction
Flourishing from training on massive data [9, 117] and high-quality human feedback [84], Large Language Models (LLMs) [84, 102, 82, 101, 46] have demonstrated remarkable abilities in instruction following and few-shot problem solving, sparking a revolution in AI field. Despite such a prosperity, LLMs remain a double-edged sword with the potential ethical risks existing before [122, 8] further amplified [110, 68, 75] or new problems emerging [8, 118], regarding particular concerns on social bias [64, 23], ethics problems [79, 36, 47], and toxicity [21, 26] in the generated content.
@@
To regulate and foster the responsible development of booming LLMs, it is necessary to assess the extent to which they conform to human values and ethics [94]. Existing approaches mostly rely on static benchmarks, e.g., RealToxicityPrompts [26] and HarmBench [74] targeting harmfulness, and ETHICS [37] and $\delta$-RoT [88] emphasizing ethical values. Nevertheless, in the era of LLMs, these datasets face the evaluation chronoeffect challenge [^1] In medicine, chrono-effectiveness refers to how a medication’s desired effects can vary with time., namely, static benchmarks are vulnerable to data leakage, hurting fair evaluation once included in training corpora [31, 54], or struggle to keep pace with fast-growing LLMs in testing difficulty, causing potential overestimation [67, 66]. As shown in Fig. 1, updated GPT models show constantly reduced toxicity on RealToxicityPrompts, yet newly constructed datasets [24, 11] reveal much more harmfulness. The latest ChatGPT also provides safer responses to the trolley problem than earlier versions.
@@
Figure 1: Illustration of evaluation chronoeffect. (a) Toxicity (the lower the better) of updated GPT versions measured on RealToxicityPrompts (upper) and toxicity of GPT-3.5-turbo (230315) on different datasets (bottom). (b) Different ChatGPT versions’ responses to the trolley problem.
For this problem, Computerized Adaptive Testing (CAT) [103] stands out as a potential solution, which utilizes Item Response Theory (IRT) [12] to model examinees’ cognitive level and adaptively selects the most appropriate next testing item from an item pool, achieving higher precision with fewer items [123]. However, traditional CAT impractically assumes the difficulty completeness of the pool [112], leaving chronoeffect unresolved.
@@
Therefore, we propose a novel framework for Generative Evolving Testing of vAlues (GETA). Instead of hypothesizing difficulty completeness of the static item pool, GETA integrates CAT and Automatic Item Generation (AIG) [29], which are theoretically unified as learning a joint distribution of item and model response with both item and value conformity as latent variables. During this process, our method jointly trains a variational IRT model and an item generator to dynamically probe the underlying moral boundaries of LLMs and adaptively generate novel and real-world test items with difficulties tailored to each examinee LLM. The generator could be iteratively optimized by collecting items beyond the boundary difficulty, enabling it to evolve alongside the LLMs’ responses. In this way, GETA avoids data leakage through on-the-fly item generation, and facilitates co-evolution of items concurrent with the model’s improvement, breaking the chronoeffect bottleneck and more accurately revealing LLMs’ real alignment with human values and ethics.
Our main contributions are: (1) To our best knowledge, we are the first to introduce psychometrics into adaptive and dynamic evaluation of LLMs’ values beyond downstream performance; (2) We propose a novel GETA framework to combine CAT with AIG and facilitate adaptive testing tailored to each LLM, mitigating evaluation chronoeffect. (3) We evaluate diverse mainstream LLMs like GPT, Gemini, LLaMA and Mistral, manifesting GETA’s superiority over previous evaluation paradigms.
@@ 2 Related works
Static Evaluation of LLMs  To reveal strengths and shortcomings of LLMs, extensive static datasets have been constructed, with emphasis shifting from in-domain NLP tasks [108, 109] to general capabilities, such as MMLU [38], AGIEval [136], BIG-bench [98] and HELM [65], covering multiple aspects, as well as specific abilities like instruction following [113, 63], domain knowledge [32, 133] and tool use [61, 130]. Besides abilities, the potential social risks [122], safety issues [17, 93] and trustworthiness [43, 110] of LLMs also become a key focus. Generally, these datasets fall into two lines: (1) Discriminative evaluation utilizing multi-choice questions or judgement [20, 37, 48, 128, 99]. (2) The generative ones carefully crafted using templates [81, 5, 80] or prompts [16, 85, 3, 18, 116] to elicit LLMs’ harmful behaviors, with BAD [129], HarmfulQA [6] and DNA [114] as typical examples. Despite widespread use, as discussed in Sec. 1, these benchmarks suffer from evaluation chronoeffect and often lack adaptability and scalability, failing to provide reliable assessment results.
@@
Dynamic Evaluation of LLMs  To compensate for the limitations of static datasets, there are growing research efforts on dynamic evaluation [56, 19]. One branch primarily follows a human-in-the-loop schema, enhancing data complexity and evaluation credibility through human interaction [72, 134, 106, 49, 10, 4], which offers greater flexibility but remains limited in scalability due to expensive human labour. Another potential direction incorporates auto-generated evaluation data through task-related structures to explicitly control test item generation, such as trees for debugging [90] and directed acyclic graphs for reasoning [137]. However, they are not suitable for our topic as it is hard to develop compositional structures of subtle human ethics. Beyond these tasks, few efforts concentrate on probing value vulnerabilities of LLMs [74, 87, 25, 40]. For instance, MasterKey [15] fine-tunes an LLM for automatic jailbreak, SAP [14] instructs LLMs to imitate human-written test prompts, and GPTFuzzer [132] leverages LLMs in a black-box fuzzing [120, 50] framework. Though such methods are unable to produce difficulty-adaptive items, this branch is poised for further exploration.
Psychometrics Based Evaluation  In psychology, psychometrics investigate the objective measurement of latent traits, e.g, intelligence [100]. Usually, a Cognitive Diagnosis Model (CDM), like Item Response Theory (IRT) [12, 124, 51], serves as the evaluation paradigm to model the probability of correct responses based on examinee ability and item difficulty. CDM is commonly combined with Computerized Adaptive Testing (CAT) [123, 103, 7] to iteratively select the next item according to the examinee’s response history, allowing direct and efficient comparison [107, 138, 139]. To reduce the high cost of item construction, Automatic Item Generation (AIG) [28] was proposed to create new items more efficiently, leveraging well-designed templates [29, 34, 33]. With the rapid development of AI, CAT has been introduced as a robust NLP metric  [73, 86] in question answering [91, 104], natural language inference [57, 58, 104] and machine translation [41, 83, 59]. More recently, this paradigm is also exploited for evaluating chatbots [95] and LLMs [140]. Nevertheless, it’s challenging to apply CAT to LLMs, since the scale invariance of IRT requires labor-intensive calibration [92] and the difficulty and scale of testing is limited due to the poor item quality by traditional AIG methods.
In spite of great progress in LLM evaluation, aforementioned limitations necessitate the integration of these methods’ advantages for better uncovering the true value boundaries of LLMs.
@@ 3 Methodology
In this section, we begin with the formalization of value evaluation and introduce static evaluation and CAT in Sec. 3.1, describe the combination of CAT and AIG in 3.2, and elaborate GETA in Sec. 3.3.
@@ 3.1 Formalization and Preliminaries
Formalization  Given a group of $m$ LLM examinees =$\mathcal{E}=\{e_{i}\}^{m}_{i=1}$ and a static dataset containing $n$ test items $\mathcal{X}=\{x_{j}\}^{n}_{j=1}$, we collect a set of responses from each LLM, $\mathcal{Y}=\{y_{i,j}\}^{m,n}_{i=1,j=1}$, where $y_{i,j}\in\{0,1\}$ indicates whether item $x_{j}$ is correctly answered by examinee $e_{i}$, aiming to accurately estimate the value conformity, $\{a_{i}\}_{i=1}^{m}$, of each LLM based on $\mathcal{X}$ and $\mathcal{Y}$. To this end, there are two main paradigms previously, namely, Static Evaluation and Adaptive Testing, as illustrated in Fig. 2.
@@
Static Evaluation  This paradigm relies on the static testing items (questions/prompts) and measures the conformity by $a_{i}=\mathbb{E}_{(x,y^{*})\sim(\mathcal{X},\mathcal{Y}^{*})}[e_{i}(y^{*}|x)]$, where $\mathcal{Y}^{*}$ denotes the ground-truth response set and $e_{i}(y^{*}|x)$ is the probability that LLM $i$ produces the correct answer [22, 1, 94]. When $\mathcal{Y}^{*}$ is unavailable, $a_{i}$ can be reformed as $\mathbb{E}_{x\sim\mathcal{X}}\mathbb{E}_{y\sim e_{i}(y|x)}[p_{\bm{\omega}}(y,x)]$ where $p_{\bm{\omega}}$ is an evaluator parameterized by $\bm{\omega}$ to tell whether the generated $y$ complies with values, like another LLM [135, 70] or fine-tuned reward models [55, 60]. However, static evaluation struggles with the chronoeffect challenge.
Figure 2: An illustration of Static Evaluation, CAT, and GETA for accessing LLMs’ value conformity.
@@
Computerized Adaptive Testing  CAT [123] was proposed in psychology to decipher the underlying latent traits of examinees more efficiently, which consists of three components. (1) An Item Response Theory (IRT) model [13] to connect the probability that $e_{i}$ correctly respond to $x_{j}$ with examinee traits and item parameters. We consider a two-parameter logistic IRT model (IRT-2PL) in this work:
$p(y_{i,j}=1|a_{i},b_{j},c_{j})=\frac{1}{1+\exp(-c_{j}(a_{i}-b_{j}))},$ (1)
where $a_{i}$ is the value conformity of $e_{i}$, $b_{j}$ and $c_{j}$ are the difficulty parameter and discrimination parameter of item $x_{j}$, respectively. (2) A calibrated item pool $\{x_{j},b_{j},c_{j}\}_{j=1}^{n}$ where the parameters of each item, $b_{j},c_{j}$, typically together with $a_{i}$, are estimated by Maximum Likelihood Estimation (MLE): $\{\hat{a}_{i}\}_{i=1}^{m},\{b_{j},c_{j}\}_{j=1}^{n}\!=\!\underset{a,b,c}{\arg\max}\prod_{i,j}p_{i,j}^{y_{i,j}}(1\!-\!p_{i,j})^{1\!-\!y_{i,j}}$, where $p_{i,j}\!=\!p(y_{i,j}=1|a_{i},b_{j},c_{j})$. (3) A selection algorithm, usually based on Fisher information [71]. After the examinee’s traits at the $t$-th testing step are measured as $\hat{a}^{t}_{i}\!=\!\underset{a_{i}}{\arg\max}\ \log{\prod_{x_{j}\in S_{t}}p_{i,j}^{y_{i,j}}(1\!-\!p_{i,j})^{1\!-\!y_{i,j}}}$, where $S_{t}\!=\!\{s_{1},...,s_{t}\}$ is the tested item sequence. The next item with maximum Fisher information, $\mathcal{F}_{\hat{a}^{t}_{i}}(x_{j})$, is selected:
$s_{t+1}\!=\!\underset{x_{j}}{\arg\max}\ \mathcal{F}(\hat{a}^{t}_{i},x_{j}),\ \mathcal{F}(a,x_{j})\!=\!{c_{j}}^{2}p(y\!=\!1|a,b_{j},c_{j})(1-p(y\!=\!1|a,b_{j},c_{j})),$ (2)
to alternatively re-estimate traits until $\hat{a}^{t}_{i}$ converges or the maximum iteration number is reached [13].

### v3

@@ Abstract
Warning: Contains harmful model outputs.
Despite significant advancements, the propensity of Large Language Models (LLMs) to generate harmful and unethical content poses critical challenges. Measuring value alignment of LLMs becomes crucial for their regulation and responsible deployment. Although numerous benchmarks have been constructed to assess social bias, toxicity, and ethical issues in LLMs, those static benchmarks suffer from evaluation chronoeffect, in which, as models rapidly evolve, existing benchmarks may leak into training data or become saturated, overestimating ever-developing LLMs. To tackle this problem, we propose GETA, a novel generative evolving testing approach based on adaptive testing methods in measurement theory. Unlike traditional adaptive testing methods that rely on a static test item pool, GETA probes the underlying moral boundaries of LLMs by dynamically generating test items tailored to model capability. GETA co-evolves with LLMs by learning a joint distribution of item difficulty and model value conformity, thus effectively addressing evaluation chronoeffect. We evaluated various popular LLMs with GETA and demonstrated that 1) GETA can dynamically create difficulty-tailored test items and 2) GETA’s evaluation results are more consistent with models’ performance on unseen OOD and i.i.d. items, laying the groundwork for future evaluation paradigms.
Machine Learning, ICML
@@ 1 Introduction
Flourishing from large-scale training (Brown et al., 2020; Wei et al., 2022a) and high-quality human feedback (Ouyang et al., 2022), Large Language Models (LLM) have demonstrated remarkable abilities in instruction following and problem-solving, sparking a revolution in the AI field. Despite such prosperity, LLMs remain a double-edged sword, with further amplified existing ethical risks (Weidinger et al., 2021b; Bommasani et al., 2022; Wang et al., 2023a; Liu et al., 2023d; McKenzie et al., 2023) and emerging new issues, particularly regarding social bias (Liang et al., 2021; Gallegos et al., 2024), ethics problems (Moor, 2006; Hendrycks et al., 2020; Jiang et al., 2022), and toxicity content (Fortuna & Nunes, 2018; Gehman et al., 2020).
@@
To ensure responsible development, it is necessary to assess to what extent LLMs conform to human values and ethics (Scherrer et al., 2023). Existing approaches mostly rely on static benchmarks, e.g., RealToxicityPrompts (Gehman et al., 2020) and HarmBench (Mazeika et al., 2024) targeting harmfulness, and ETHICS (Hendrycks et al., 2021a) and $\delta$-RoT (Rao et al., 2023) emphasizing ethical values. However, these datasets face the evaluation chronoeffect challenge, namely, i) static benchmarks are vulnerable to data leakage, hurting fair evaluation once leaking into training corpora (Golchin & Surdeanu, 2023; Kocoń et al., 2023), and ii) quick saturation with fast-growing LLMs in terms of testing difficulty, causing potential overestimation due to ceiling effect (Liu et al., 2023a, b). As shown in Fig. 1, although updated versions of GPT models show constantly reduced toxicity on RealToxicityPrompts, newly constructed datasets (Ganguli et al., 2022; Cui et al., 2023) reveal much more harmfulness.
@@
Figure 1: Illustration of evaluation chronoeffect. (a) Toxicity ($\downarrow$) of updated GPT versions measured on RealToxicityPrompts (upper) and toxicity of GPT-3.5-turbo (230315) on different datasets (bottom). (b) Different ChatGPT versions’ answers to the trolley problem.
Similar problems arise in psychometrics, where new tests are necessary as human skills advance, but their effectiveness diminishes once they become the focus of learning. For this problem, Computerized Adaptive Testin (CAT; van der Linden & Glas, 2010) stands out as a potential solution in psychometrics, which utilizes Item Response Theory (IRT; De Ayala, 2013) to model examinees’ capability and adaptively selects the most appropriate test item for each examinee from an static item pool, aiming at using fewest items (Weiss & Kingsbury, 1984) to discriminate different levels of capabilities. Although such methods effectively increase test longevity, they are constrained by the static item pool, which fails to adapt to the rapidly growing model capabilities and prevent data leakage, leaving chronoeffect in LLM evaluation unresolved.
@@
In this paper, we propose a novel framework for Generative Evolving Testing of vAlues (GETA) to dynamically evaluate LLMs’ value alignment. Without relying on the static item pool, GETA integrates CAT with Automatic Item Generation (AIG; Gierl et al., 2012), which generates new test items of varying difficulty. Our method jointly trains a Variational IRT (VIRT) model and an item generator to dynamically probe the underlying moral boundaries of LLMs and adaptively generate novel test items with difficulties tailored to each examinee LLM. The generator could be iteratively optimized by collecting items beyond the difficulty boundary, allowing it to evolve alongside the LLMs’ responses. In this way, GETA avoids data leakage through on-the-fly item generation and co-evolves with examinee LLMs’ improvement, addressing the chronoeffect challenge.
Our main contributions are: (1) We introduce psychometrics methods into adaptive and dynamic evaluation of LLMs’ value alignment ; (2) We propose GETA, a novel framework integrating CAT with AIG, to address evaluation chronoeffect. (3) ate GETA’s benefits over previous evaluation paradigms, including static and adaptive methods, by evaluating mainstream LLM’s value conformity.
@@ 2 Related works
Figure 2: An illustration of Static Evaluation, CAT, and GETA: (1) The static evaluation method directly applies existing benchmarks to LLMs and computes the metrics. (2) The CAT method first calibrates the static data with human responses at scale to create an item pool and fit an IRT model. Then, at each testing step, the best-fitting item is selected from the pool for each examinee LLM, and the ability of each examinee LLM is estimated based on its response history, both using the IRT model. (3) GETA jointly trains a variational IRT model and an item generator powered by a pre-trained language model with the calibrated items, thereby adaptively generating difficulty-tailored test items instead of relying on static items, addressing the chronoeffect challenge. A running example of GETA is shown on the right.
Static Evaluation of LLMs  Extensive static datasets have been constructed to evaluate LLM’s capability, from in-domain NLP tasks (Wang et al., 2018, 2019) to general capabilities, such as MMLU (Hendrycks et al., 2021b), BIG-bench (Srivastava et al., 2023), and HELM (Liang et al., 2023). Recently, social risks (Weidinger et al., 2021b), safety issues (Dong et al., 2024; Röttger et al., 2024) and trustworthiness (Huang et al., 2023; Wang et al., 2023a) of LLMs have become a key evaluation focus. However, static benchmarks often suffer from the chronoeffect challenge. The common issues such as test data leakage could invalidate benchmark results; the rapidly advancing LLMs can quickly saturate a static benchmark, making it unable to discriminate models with various levels of capability.
@@
Dynamic Evaluation of LLMs  There are growing research efforts on dynamic evaluation (Krause et al., 2018; Fan et al., 2024) to address test data leakage. One branch primarily follows a human-in-the-loop schema to enhance data complexity and evaluation credibility (Ma et al., 2021; Zellers et al., 2021; Vidgen et al., 2021; Kiela et al., 2021; Collins et al., 2023), which offers greater flexibility but remains limited in scalability due to expensive human labour. Another direction explicitly guides test item generation through task-related data structures, such as trees for debugging (Ribeiro & Lundberg, 2022) and directed acyclic graphs for reasoning (Zhu et al., 2024). However, those methods are not suitable for value alignment evaluation as no compositional structures can be developed for nuanced human ethics. For value alignment, a few efforts concentrate on probing value vulnerabilities of LLMs (Mazeika et al., 2024; Radharapu et al., 2023; Ge et al., 2023; Hong et al., 2024). For instance, MasterKey (Deng et al., 2023b) fine-tunes an LLM for automatic jailbreak, SAP (Deng et al., 2023a) instructs LLMs to imitate human-written test prompts, and GPTFuzzer (Yu et al., 2023) leverages LLMs in a black-box fuzzing (Wei et al., 2018; Kim et al., 2020) framework. However, those methods were primarily designed for identifying vulnerabilities rather than systematically evaluating value alignment, struggling to ensure comprehensive assessment and adaptability to evolving ethical considerations.
Adaptive Testing in Psychometrics  In psychometrics, Computerized Adaptive Testing (CAT; Weiss & Kingsbury, 1984; van der Linden & Glas, 2010; Bi et al., 2021) is commonly used to create efficient personalized tests. CAT adaptively selects test items from the pool based on an Item Response Theory (IRT) model that models the probability of correct responses based on the examinee’s ability and test item parameters  (De Ayala, 2013; Wu et al., 2020; Kim et al., 2023). Such a method could minimize the number of test items required to estimate the capability of interests while maximizing the test’s discriminative power. CAT has been introduced into AI evaluation, such as evaluations of question answering (Rodriguez et al., 2021; Vania et al., 2021), machine translation (Hopkins & May, 2013; Otani et al., 2016; Lalor et al., 2019), as well as chatbots and LLMs (Sedoc & Ungar, 2020; Zhuang et al., 2023, 2024; Polo et al., 2024; Lalor et al., 2024). Although CAT leads to efficient, dynamic benchmarks, they are constrained by the quality and size of the static item pool (e.g., test dataset) and are still vulnerable to data leakage and difficulty saturation.
@@ 3.1 Formalization and Preliminaries
Formalization  Given a group of $m$ examinee LLMs =$\mathcal{E}\!=\!\{e_{i}\}^{m}_{i=1}$ and a static dataset containing $n$ test items $\mathcal{X}\!=\!\{x_{j}\}^{n}_{j=1}$, we collect responses from each LLM, denoted as =$\mathcal{R}\!=\!\{r_{i,j}\}^{m,n}_{i=1,j=1}$, where $r_{i,j}$ represents the response of examinee $e_{i}$ to item $x_{j}$. The correctness of $\mathcal{R}$ is defined as $\mathcal{Y}\!=\!\{y_{i,j}\}^{m,n}_{i=1,j=1}$ with $y_{i,j}\!\in\!\{0,1\}$ indicating whether $r_{i,j}$ aligns with human values. $(\mathcal{X},\mathcal{Y})$ is then used to estimate the value conformity $\{a_{i}\}_{i=1}^{m}$ of each LLM. To this end, two primary paradigms have been established previously: Static Evaluation and Adaptive Testing, as illustrated in Fig. 2.
@@
Static Evaluation (SE)  This paradigm relies on the static test questions and calculates value conformity as $a_{i}\!=\!\mathbb{E}_{(x,r^{*})\sim(\mathcal{X},\mathcal{R}^{*})}[e_{i}(r^{*}|x)]$, where $\mathcal{R}^{*}$ denotes the set of ground-truth response $r^{*}$ and $e_{i}(r^{*}|x)$ is the probability that LLM $e_{i}$ produces the correct answer (Fraser et al., 2022; Arora et al., 2023; Scherrer et al., 2023). When $\mathcal{R}^{*}$ is unavailable, $a_{i}$ can be reformulated as $\mathbb{E}_{y\sim\mathcal{Y}}(y)$ where $y$ is determined by an evaluator designed to assess whether the response $r$ complies with specified values, such as another LLM (Zeng et al., 2023; Liu et al., 2023c) or fine-tuned reward models (Köpf et al., 2023; Lambert et al., 2024). However, SE struggles with the chronoeffect challenge.
@@
Computerized Adaptive Testing  CAT (Weiss & Kingsbury, 1984) was proposed to efficiently decipher the latent psychological traits of examinees, consisting primarily of three components: (1) An IRT model (de Ayala, 2022) that connects the probability of $e_{i}$ correctly responding to $x_{j}$ with examinee ability $a_{i}$ and item parameters $b_{j}$, $c_{j}$. Here, $b_{j}$ is item difficulty, indicating the item’s position on the difficulty scale. $c_{i}$ is item discrimination, which describes how sharply the success probability changes with ability $a_{i}$. We adopt a two-parameter logistic IRT model (IRT-2PL):
$p(y_{i,j}=1|a_{i},b_{j},c_{j})=\frac{1}{1+\exp(-c_{j}(a_{i}-b_{j}))}.$ (1)
(2) A calibrated item pool $\{x_{j},b_{j},c_{j}\}_{j=1}^{n}$, where the item parameters $b_{j},c_{j}$ and the examinee ability $a_{i}$ are estimated via Maximum Likelihood Estimation (MLE): $\{\hat{a}_{i}\}_{i=1}^{m},\{b_{j},c_{j}\}_{j=1}^{n}\!=\!\underset{a,b,c}{\arg\max}\prod_{i,j}p_{i,j}^{y_{i,j}}(1\!-\!p_{i,j})^{(1\!-\!y_{i,j})}$, where $p_{i,j}\!=\!p(y_{i,j}\!=\!1|a_{i},b_{j},c_{j})$, based on a large human response set. (3) A selection algorithm to select the next appropriate item for testing. At the $t$-th step, the examinee ability is measured as $\hat{a}^{t}_{i}\!=\!\underset{a_{i}}{\arg\max}\ \log{\prod_{x_{j}\in S_{i}^{t}}p_{i,j}^{y_{i,j}}(1\!-\!p_{i,j})^{(1\!-\!y_{i,j})}}$, where $S_{i}^{t}\!=\!\{s_{i}^{1},...,s_{i}^{t}\}$ is the tested item sequence. Then the next item is selected by maximizing the Fisher information $\mathcal{F}_{\hat{a}^{t}_{i}}$ (Ly et al., 2017): $s_{i}^{t+1}\!=\!\underset{x_{j}\in\mathcal{X}}{\arg\max}\ \mathcal{F}_{\hat{a}^{t}_{i}}(b_{j},c_{j}),\ \mathcal{F}_{a_{i}}(b_{j},c_{j})\!=\!c_{j}^{2}\cdot p_{i,j}(1-p_{i,j}).$ CAT iteratively updates $\hat{a}^{t}_{i}$ and adaptively selects $s^{t}_{i}$ until a certain termination criterion is met. While CAT requires minimal data for testing, its static item pool may lead to overestimation due to insufficiently challenging items. A detailed description of CAT is in Appendix. C.1.
Algorithm 1 GETA Algorithm
  Input: $\mathcal{E}$, $q_{\bm{\theta}}$, $q_{\bm{\phi}}$, $p_{\bm{\omega}}$, $\{(x^{0},d^{0})\}$, $T$, $k_{1}$, $k_{2}$, $\delta_{1}$, $\delta_{2}$ and $\mathcal{D}=\emptyset$
  Output: $\{\hat{a}^{T}_{i}\}^{m}_{i=1}$ and the evolved $p_{\bm{\omega}}(x|d)$
  for $i=1,2,...,m$ do
     Sample $y_{i}^{0}\!\sim\!e_{i}(y|x^{0})\text{for each}\ x^{0}$
     Calculate $\hat{a}_{i}^{0}$ with $q_{\bm{\theta}}$, $S_{i}^{0}\!=\!\{(x^{0},d^{0},y_{i}^{0})\}$
  end for
  for $t=1,2,...,T$ do
     for $i=1,2,...,m$ do
        Calculate $d^{*}$ for $e_{i}$ with Eq. (5),
        Sample $x_{j}^{t}$ with Eq. (6), $j\!=\!1$ to $k_{1}$
        Sample $y_{i,j}^{t}\!\sim\!e_{i}(y|x_{j}^{t})\text{for each}\ x_{j}^{t},e_{i}$
        Calculate $\hat{d}_{j}^{t}$ with $q_{\bm{\phi}}$
        for $j=1,2,...,k_{1}$ do
           if $|\hat{d}_{j}^{t}-d^{*}|<\delta_{1}$ then
              $S_{i}^{t-1}\leftarrow S_{i}^{t-1}\cup\{(x_{j}^{t},\hat{d}_{j}^{t},y_{i,j}^{t})\}$
           else if $|\hat{d}_{j}^{t}-d^{*}|>\delta_{2}$ then
              $\mathcal{D}\leftarrow\mathcal{D}\cup\{(x_{j}^{t},\hat{d}_{j}^{t})\}$
           end if
        end for
        $S_{i}^{t}\leftarrow S_{i}^{t-1}$, Calculate $\hat{a}_{i}^{t}$ with $q_{\bm{\theta}}$
     end for
     if $|\mathcal{D}|\geq t*k_{2}$ then
        Optimize $\bm{\omega}$ on $\mathcal{D}$
     end if
  end for